Problem 2: Find the nodal voltages and the current I. 30 40 V 60
Answer #1

į Jy nodal analysis at node asb yra t 3va -21 +6vb tar = o 7V +8V6 =21 - 0) Va-vb = -3 - big va = -3 + - (119) From ef? og 01No of eq? (131) Vo in but value va = 16-3 = -0.7V I a 7 - Va 4 a 7 - (-03) ។ = 7 + 0.7 4 i I :18A va = -0.27 Vo = 2.8V I = 1.